celticcat:

The differential equation below models the temperature of a 93°C cup of coffee in a 19°C room, where it is known that the coffee cools at a rate of 1°C per minute when its temperature is 70°C. Solve the differential equation to find an expression for the temperature of the coffee at time t. dy dt = − 1 51 (y − 19). Not 10%% sure about this one I got a general solutioin y = c e^(-1.51t) + 19 but I cant see how to find the constant c.

justjm:

The differential is \(\frac{dT}{dt} = \frac{-1}{51}(T-19) \) right?

celticcat:

Ill just check that ..

celticcat:

I took it to mean -1.51 but its -1/51.

justjm:

Yeah so I got \(y=Ce^{-\frac{1}{51}t} +19\) Now to find C, you can plug in the initial temperature (93, 0), setting y=93 and t=0.

celticcat:

OK THanks.

justjm:

np, lmk what you get for C. After that's done to find the temp at 70°, you'll just set y=70 and solve for t

celticcat:

Sorry I got called away. I get C to be 74 tight?

justjm:

Yes

celticcat:

19 secionds for time when its 70 degrees?

justjm:

18.984≈19 seconds, you're correct

celticcat:

Thanks for your help!!

justjm:

np